Solve in series 2x(1-x)d^2y/dx^2+(1-x)dy/dx+3y=0


1
Expert's answer
2021-06-18T12:21:56-0400

Solution

Let

"y(x)=\sum_{n=0}^{\infty}{a_nx^n}"

"\frac{dy}{dx}=\sum_{n=1}^{\infty}{na_nx^{n-1}}=\sum_{n=0}^{\infty}{(n+1)a_{n+1}x^n}"

"\frac{d^2y}{dx^2}=\sum_{n=0}^{\infty}{n(n+1)a_{n+1}x^{n-1}}=\sum_{n=0}^{\infty}{(n+1)(n+2)a_{n+2}x^n}"

"(1-x)\frac{dy}{dx}=a_1+\sum_{n=1}^{\infty}[(n+1)a_{n+1}-na_n]x^n"

"2x(1-x)\frac{d^2y}{dx^2}=4a_2x+2\sum_{n=0}^{\infty}{[(n+1)na_{n+1}-(n-1)na_n]x^n}"

Substitution into equation:

"4a_2x+2\sum_{n=2}^{\infty}{\left[n\left(n+1\right)a_{n+1}-\left(n-1\right)na_n\right]x^n}+a_1+\left[2a_2-a_1\right]x+\sum_{n=2}^{\infty}{\left[\left(n+1\right)a_{n+1}-na_n\right]x^n}+3a_0+3a_1x+3\sum_{n=2}^{\infty}{a_nx^n}=0"

Coefficient near xn are equal to zero.

x0 : 3a0+a1=0 => a1=-3a0 

x1 :4a2+2a2-a1+3a1=0 => 6a2+2a1=0 => a2=-a1/3= a0 

xn (n>1):2n(n+1)an+1-2(n-1)nan+(n+1)an+1-nan+3an=0 => (2n+1)(n+1)an+1 +(3+n-2n2)an=0 =>

an+1 =(2n2-n-3)an/[(2n+1)(n+1)]=(2n-3)an/(2n+1)

a3 =a2/5= a0/5, a4 =3a3/7= 3a0/(5*7), a5 =5a4/9=5*3a0/(5*7*9), a6 =7a5/11=7*5*3a0/(5*7*9*11), …,

an+1 =3a0/[(2n+1)(2n-1)]

General formula for an: an =3a0/[(2n-1)(2n-3)], n>0

Therefore solution is

"y(x)=3a_0\sum_{n=0}^{\infty}\frac{x^n}{(2n-1)(2n-3)}"